# Scalable and shallow quantum circuits encoding probability distributions informed by asymptotic entanglement analysis

### v2

@@ Abstract
1 Introduction
The first step of many quantum algorithms is state preparation, encoding the classical input data in e.g. the amplitudes of a quantum state. For a general input, the depth of the quantum circuit preparing a corresponding quantum state may scale exponentially [1], making this key step intractable on real quantum devices. To alleviate this problem different strategies have been proposed, including general methods based on Quantum Generative Adversarial Networks (QGANs)[2], Quantum Fourier Transform (QFT) [3, 4, 5], or handcrafted approaches [6], but all of these have shortcomings in terms of scaling, circuit complexity, trainability or generality, limiting their usefulness in practice. Fortunately, for important classes of data accurate approximate and only polynomially deep circuits can be systematically constructed. This is based on two properties: First, the representability of the input data as a matrix product state (MPS)[7, 8] [more generally: a tensor network (TN)] of small bond dimension. Second, the ability to convert a small bond dimension MPS into a quantum circuit with gate count linear in the number of qubits [9, 10, 11, 12, 13, 14].
@@ 1 Introduction
For the first question, it is known that while a generic function may require an MPS approximation of an exponential rank (bond dimension), polynomials are represented exactly with small bond MPS [15]. Further, results for ranks and error bounds of low-rank MPS approximations exist for smooth functions [16, 17, 18, 19, 20], some of which were used in the previous works on distribution encoding [21, 22, 23, 24, 25]. Here, we significantly strengthen these analyses by deriving results on the precise structure of the entanglement in the MPS approximation in terms of entanglement spectra and purities. In particular we show universal scaling relations of these entanglement measures for smooth functions, and their breakdown in the absence of smoothness. These allow us to derive analytically functionals controlling e.g. the fidelity of a single layer approximate encoding circuits.
@@
Finally, motivated by their role in quantum finance routines [26, 27], we deploy our algorithm in the special case of encoding probability distributions, particularly the $\alpha$-stable ones, i.a. the Lévy distribution, but also other examples. We test our methods by executing the circuits on real and simulated IBM quantum devices. In particular we execute and sample from the quantum circuits generated by our method on the ibm_torino QPU on up to 25 qubits, with the generated distributions passing standard statistical tests (e.g. Kolmogorov-Smirnov). We also execute very large circuits for up to 64 qubits on the ibm_marrakesh device, which we generate using a scalable compute- and memory-efficient Tensor Cross Interpolation [28] extension of our algorithm, obtaining qualitative agreement with the exact distributions even for these large, utility-scale problems.
We particularly focus on encoding the Lévy and $\alpha$-stable distributions, well-suited for modeling heavy-tailed or skewed data and volatility clustering, commonly observed in financial markets. Unlike Gaussian distributions, which fail to capture the frequency of extreme market events, they more effectively model the fat tails of financial data. These properties makes them invaluable in risk assessment, portfolio management, and derivative pricing [29, 30]. Specifically, $\alpha$-stable distributions are a natural choice whenever stability under addition is needed, to ensure consistency when aggregating risks or returns from a portfolio [31, 32] or predicting risk on different time horizons, as in Target or Goal-oriented investment strategies [33, 34]. In Option Pricing, generalizations of the Black-Scholes model incorporating Lévy distributions (e.g. the CGMY process) provide a better fit for market data by accounting for jumps and heavy tails [30]. In Risk Management, in stress testing and value-at-risk (VaR) calculations, $\alpha$-stable models help to quantify risks under non-Gaussian scenarios, reflecting market crashes and high-volatility periods [35, 36]. In High-Frequency Trading Lévy processes are used in high-frequency data analysis, where non-Gaussian features like skewness and kurtosis are prominent [31, 37]. Finally, for multiple asset portfolios the total portfolio risk can be calculated from its constituents: the sum of the risks retains the Lévy distribution properties, enabling consistent modeling across scales [35, 29]. We show that such distributions can efficiently be loaded on QPUs.
The paper is structured as follows: In Sec.2 we introduce the problem and the necessary background. In Sec.3 we state our theoretical results, numerical experiments, and discuss their implications. In Sec.4 we give an improved circuit construction algorithm based on them. In Sec.5 we benchmark the results on real and simulated quantum hardware, and in Sec.6 we discuss future directions. Appendices give mathematical details of the proofs, and further algorithmic details.
2 Preliminaries: MPS approximation of functions
@@
$c_{\sigma_{1}\sigma_{2}\dots\sigma_{N}}=\Gamma^{\sigma_{1}}_{1,l_{1}}\Lambda_{1,l_{1}}\Gamma^{\sigma_{2}}_{2,l_{1}l_{2}}\Lambda_{2,l_{2}}\dots\Gamma^{\sigma_{N-1}}_{N-1,l_{N-2},l_{N-1}}\Lambda_{N-1,l_{N-1}}\Gamma^{\sigma_{N}}_{N,l_{N-1}},$ (5)
@@
In what follows the MPS approximation of the function will be converted to a shallow quantum circuit preparing the encoded function on a quantum device (Fig.1d), using our improved algorithm (see Sec.4) extending the methods of Refs.[21, 22, 23, 24, 25].
3 Entanglement in tensorial approximants of smooth functions
3.1 General results
@@
$\rho_{k}=|\sin{\pi x}\rangle\left(\frac{1}{2}+\frac{2^{k}}{4\pi}\sin(2\pi/2^{k})\right)\langle\sin{\pi x}|+|\cos{\pi x}\rangle\left(\frac{1}{2}-\frac{2^{k}}{4\pi}\sin\left(2\pi/2^{k}\right)\right)\langle\cos{\pi x}|+\\ +\frac{2^{k}}{2\pi}\sin^{2}{(\pi/2^{k})}(|\cos{\pi x}\rangle\langle\sin{\pi x}|+|\sin{\pi x}\rangle\langle\cos{\pi x}|)$ (12)
@@
Figure 2: Comparison of the analytical (see Theorem 1 and Corollary 1) and numerical SVD results on the dependence of the entanglement spectra $\Lambda_{k,1}$ and $\Lambda_{k,2}$ on bond position $k$. a) $\Lambda_{k,1}$ for the Lévy distribution with $c=0.15$ and $c=4$. b) $\Lambda_{k,1}$ for the normal distribution with $\mu=0.5$ for $\sigma=0.125$ and $\sigma=0.5$. c) $\Lambda_{k,2}$ for the Lévy distribution with $c=0.15$ and $c=4$. d) $\Lambda_{k,2}$ for the normal distribution with $\mu=0.5$ for $\sigma=0.125$ and $\sigma=0.5$. Note the excellent agreement and universal behaviour for large $k$, and non-universal features for small $k$, as discussed in the text. Observe also the larger values of entanglement for the the less localized functions (with larger parameters $c$ and $\sigma$).
The above results are based on function smoothness, and characterize the asymptotic behavior for qubits describing the smallest spatial scales of function’s variations (as per Eq.1). The entanglement spectra for the leading qubits, corresponding to the large scales, are non-universal, but we observe a strong dependence on the localization properties of the function. For example, a maximally localized delta-peaked distribution is described by an entirely unentangled product state. Similarly, as described in more detail below, exponentially decaying functions are less entangled than the power-law tailed ones.
@@
3.2 Encoding probability distributions
@@
Finally, in Figs.2c,d we compute and compare the third entanglement eigenvalue $\Lambda_{k,2}$ for Lévy and normal distributions, respectively. As expected, our analytical result matches the numerical data, which indeed behave as $O(1/4^{k})$. As in the previous cases, a discrepancy with analytic results appears only on the first bonds corresponding to the large scales. Note also, that the entanglement is larger in the less localized functions, i.e. Lévy distributions with a larger parameter $c$ and normal distributions with a larger $\sigma$.
@@
3.3 Encoding dependence on the localization and support of the function
@@
The above argument is borne out by numerical experiments. In Figs.3a,b we plot the logarithm of the second entanglement eigenvalue $\Lambda_{k,1}$ on bond $k$ between qubits $k$ and $k+1$, for the normal and Lévy distributions, respectively. The large-$k$ tails are universal and exponentially decaying, as derived in Sec.3.1. The low-$k$ behaviour, however, shows indeed a super-exponential decay. Moreover, there is a clear distinction between the exponential localization of the normal distribution and the weaker polynomial one of the Lévy distribution: for the former, in Fig.3a, the large scale (small $k$) drop of entanglement is almost instantaneous, and also the universal asymptotic regime of exponential decay for lower scales (large $k$) is clearly visible already for a heavily truncated function support of $[0,2^{4}]$. Note that the maximum is achieved between $k=3,4$, consistent with the condition $k\approx\log_{2}{(L/2\sigma)}$ derived in Sec.3.1. For the Lévy case the small-$k$ decay of entanglement is much slower (though still super-exponential), and correspondingly the onset of the universal regime of entanglement decay at small scales is pushed to much larger values of $k$: we required a support truncated to a very large interval $[0,2^{19}]$ in Fig.3b to clearly demonstrate both the low- and high-$k$ regimes.
@@
4 Constructing the encoding circuit using the MPS representation
@@
5 Numerical tests of the encoding circuits on simulators and QPUs
@@
In Fig. 6a we show the infidelity of a 1-layer approximating quantum circuit as a function of the truncated support size $L$, assuming a fixed discretization size (thus the qubit number is $N=10+\lceil\log_{2}L\rceil$). Here we consider the example of a normal distribution with $\sigma=2$. The infidelity is computed on an MPS simulator with a very large bond dimension (i.e. essentially exact simulation). Overall, the infidelity is excellent already for this extremely shallow circuit, no higher than $10^{-3}$. Note also that for smaller $L$ it perfectly follows the analytical estimate from Eq.(17). For larger supports this simple approximation is not accurate: as discussed in Sec.3.1 the smoothness-controlled regime is valid from bond $k=\log_{2}(L/2\sigma)$ upwards, and below the entanglement is localization controlled (and, for this exponentially localized function, much smaller, see Fig.3a). For large $L$ thus Eq.(17) overestimates the infidelity, which can be corrected by only summing contributions from bonds $k\geq log_{2}(L/2\sigma)$, i.e. the ones for which the asymptotic scaling is valid. This corrected estimate agrees perfectly with the numerical results for all $L$.
In Fig.5 we examine the influence of using multiple disentangling V-layers in the quantum circuit construction, and the sensitivity to the V-layer origin position, on the example of the Lévy distribution. We observe that the fidelity can be systematically improved by adding additional disentangling layers (note, though, that in the presence of noise this is not true, as the increased circuit depth may result in a lower fidelity in practice). Furthermore, for more than a single layer we also see a strong dependence on the layer origin, as expected from the entanglement analysis in Sec.3.1: the effect is comparable in size to adding another disentangling layer, and hence the freedom to chose it is very important in practice. Note also that the minimal circuit depth is achieved with the origin on the middle qubits, further emphasizing the usefulness of this strategy.
@@
A more rigorous test of performance is obtained by computing statistical metrics for the samples drawn from the encoded distributions. We use the Kullback-Leibler (KL) divergence between the encoded and ideal distributions, and the Kolmogorov-Smirnov (KS) test. The results are reported in Tables 2 and 2. For these tests we use a 2-layer encoding circuit constructed by our complete algorithm as described in Sec.4, including truncation of gates in V-layers set by the parameter $\varepsilon_{trunc}$. We report the depth and amount of CNOT gates of the circuit, which are computed in the RX, RY, RZ and CNOT basis (the depth includes all one-qubit gates). To collect the data we varied the interval $L$ and scale factors of the distributions, and we report the mean value and the standard deviation. The constructed distributions mostly pass the statistical tests with a wide margin, but we note that for very small amount of qubits KS test may occasionally fail (p-value$<0.05$) due to extreme discretization of the distribution, as the test is performed against the ideal truncated cumulative density function which is continuous. In Table 2 we set the hyperparameter $\varepsilon_{trunc}$ to $10^{-3}$, while in Table 2 to $\varepsilon_{trunc}$ to $10^{-4}$. Changing this value represents an accuracy-depth trade-off, which will be important on noisy QPUs.
@@ Acknowledgements
We thank Vladyslav Los for his technical support in executing the quantum circuits on the IBM quantum hardware.
@@ A.1 Proof of Theorem 1
$p_{k}=\frac{1}{4^{k}}\sum_{\mathbf{u_{k},u^{\prime}_{k}}}\int_{0}^{1}\int_{0}^{1}f(\mathbf{u_{k}}+v/2^{k})f^{*}(\mathbf{u_{k}}+v^{\prime}/2^{k})f(\mathbf{u^{\prime}_{k}}+v^{\prime}/2^{k})f^{*}(\mathbf{u^{\prime}_{k}}+v/2^{k})\rm{d}v\rm{d}v^{\prime}.$ (20)
@@
$\sum_{u_{n}}\frac{f(u_{n})}{2^{k}}=\int_{0}^{1}f(u)\rm{d}u+\frac{f(0)-f(1)}{2^{k+1}}+B_{2}\frac{f^{\prime}(1)-f^{\prime}(0)}{2^{2k+1}}+O(1/2^{3k}),$ (21)
@@
$p_{k}=\int_{0}^{1}\int_{0}^{1}\left(\int_{0}^{1}f(u+v/2^{k})f^{*}(u+v^{\prime}/2^{k})\rm{d}u+b_{1}(v,v^{\prime})\right)\\ \left(\int_{0}^{1}f^{*}(u^{\prime}+v/2^{k})f(u^{\prime}+v^{\prime}/2^{k})\rm{d}u^{\prime}+b_{2}(v,v^{\prime})\right)\rm{d}v\rm{d}v^{\prime},$ (22)
@@
$(h_{2,0}+h_{0,2})/6+h_{1,1}/4-(|f(1)^{2}|^{\prime}-|f^{2}(0)|^{\prime})/6=\\ =(h_{2,0}+2h_{1,1}+h_{0,2})/6-h_{1,1}/12-(|f(1)^{2}|^{\prime}-|f^{2}(0)|^{\prime})/6=-h_{1,1}/12$ (27)

### v3

# Entanglement scaling in matrix product state representation of smooth functions and their shallow quantum circuit approximations

@@ Abstract
I Introduction
**footnotetext: These authors contributed equally to this work
The first step of many quantum algorithms is state preparation, encoding the classical input data in e.g. the amplitudes of a quantum state. For a general input, the depth of the quantum circuit preparing a corresponding quantum state may scale exponentially [1], making this key step intractable on real quantum devices. To alleviate this problem different strategies have been proposed, including general methods based on Quantum Generative Adversarial Networks (QGANs)[2], Quantum Fourier Transform (QFT) [3, 4, 5], or handcrafted approaches [6], but all of these have shortcomings in terms of scaling, circuit complexity, trainability or generality, limiting their usefulness in practice. Fortunately, for important classes of data accurate approximate and only linearly deep circuits can be systematically constructed. This is based on two properties: First, the representability of the input data as a matrix product state (MPS)[7, 8] [more generally: a tensor network (TN)] of small bond dimension. Second, the ability to convert a small bond dimension MPS into a quantum circuit with gate count linear in the number of qubits [9, 10, 11, 12, 13, 14].
@@ I Introduction
For the first question, it is known that while a generic function may require an MPS approximation of an exponential rank (bond dimension), polynomials are represented exactly with small bond MPS [15]. Further, results for ranks and error bounds of low-rank MPS approximations exist for smooth functions [16, 17, 18, 19, 20], some of which were used in the previous works on distribution encoding [21, 22, 23, 24, 25, 26]. Here, we significantly strengthen these analyses by deriving results on the precise structure of the entanglement in the MPS approximation in terms of entanglement spectra and purities. In particular we show universal scaling relations of these entanglement measures for smooth functions, and their breakdown in the absence of smoothness. These allow us to derive analytically functionals controlling e.g. the fidelity of a single layer approximate encoding circuits.
@@
Finally, motivated by their role in quantum finance routines [27, 28], we deploy our algorithm in the special case of encoding probability distributions, particularly the $\alpha$-stable ones, i.a. the Lévy distribution, which are used to effectively model the fat tailed or skewed data in risk assessement, portfolio management and derivative pricing [29, 30, 31, 32, 33, 34, 30, 35, 36, 37]. We test our methods by executing the circuits on real and simulated IBM quantum devices. In particular we execute and sample from the quantum circuits generated by our method on the ibm_torino QPU on up to 25 qubits, with the generated distributions passing standard statistical tests (e.g. Kolmogorov-Smirnov). We also execute very large circuits for up to 64 qubits on the ibm_marrakesh device, which we generate using a scalable compute- and memory-efficient Tensor Cross Interpolation [38] extension of our algorithm, obtaining qualitative agreement with the exact distributions even for these large, utility-scale problems.
The paper is structured as follows: In Sec.II we introduce the problem and the necessary background. In Sec.III we state our theoretical results, numerical experiments, and discuss their implications. In Sec.IV we give an improved circuit construction algorithm based on them. In Sec.V we benchmark the results on real and simulated quantum hardware, and in Sec.VI we discuss future directions. Appendices give mathematical details of the proofs, and further algorithmic details.
II Preliminaries: MPS approximation of functions
@@
$c_{\sigma_{1}\sigma_{2}\dots\sigma_{N}}=$ (5)
$\Gamma^{\sigma_{1}}_{1,l_{1}}\Lambda_{1,l_{1}}\Gamma^{\sigma_{2}}_{2,l_{1}l_{2}}\Lambda_{2,l_{2}}\dots\Gamma^{\sigma_{N-1}}_{N-1,l_{N-2},l_{N-1}}\Lambda_{N-1,l_{N-1}}\Gamma^{\sigma_{N}}_{N,l_{N-1}},$
@@
In what follows the MPS approximation of the function will be converted to a shallow quantum circuit preparing the encoded function on a quantum device (Fig.1d), using our improved algorithm (see Sec.IV) extending the methods of Refs.[21, 22, 23, 25, 26] (see also [47]).
III Entanglement in tensorial approximants of smooth functions
III.1 General results
@@
Figure 2: Comparison of the analytical (see Theorem 1 and Corollary 1) and numerical SVD results on the dependence of the entanglement spectra $\Lambda_{k,1}$ and $\Lambda_{k,2}$ on bond position $k$. a) $\Lambda_{k,1}$ for the Lévy distribution with $c=0.15$ and $c=4$. b) $\Lambda_{k,1}$ for the normal distribution with $\mu=0.5$ for $\sigma=0.125$ and $\sigma=0.5$. c) $\Lambda_{k,2}$ for the Lévy distribution with $c=0.15$ and $c=4$. d) $\Lambda_{k,2}$ for the normal distribution with $\mu=0.5$ for $\sigma=0.125$ and $\sigma=0.5$. Note the excellent agreement and universal behaviour for large $k$, and non-universal features for small $k$, as discussed in the text.
@@
$\rho_{k}=|\sin{\pi x}\rangle\left(\frac{1}{2}+\frac{2^{k}}{4\pi}\sin(2\pi/2^{k})\right)\langle\sin{\pi x}|+\\ +|\cos{\pi x}\rangle\left(\frac{1}{2}-\frac{2^{k}}{4\pi}\sin\left(2\pi/2^{k}\right)\right)\langle\cos{\pi x}|+\\ +\frac{2^{k}}{2\pi}\sin^{2}{(\pi/2^{k})}(|\cos{\pi x}\rangle\langle\sin{\pi x}|+|\sin{\pi x}\rangle\langle\cos{\pi x}|)$ (12)
@@
The above results are based on function smoothness, and characterize the asymptotic behavior for qubits describing the smallest spatial scales of function’s variations (as per Eq.1). The entanglement spectra for the leading qubits, corresponding to the large scales, are non-universal, but we observe a strong dependence on the localization properties of the function. For example, a maximally localized delta-peaked distribution is described by an entirely unentangled product state. Similarly, as described in detail below, exponentially decaying functions are less entangled than the power-law tailed ones.
Let us compare our results with those previously obtained in the literature. In Ref. [18] a bound for the purity was obtained in the following form: $p_{k}\geq 1-4\max|f(x)f^{\prime}(x)|/2^{k}$ for a positive $f$. In contrast, our analytic expansions for purities show that the first non-vanishing term scales $1/4^{k}$ for a general $f$, i.e. an exponential improvement in the bound. In Ref.[16] an Ansatz for entanglement spectra of the form $\sqrt{\rho_{kn}}=\Lambda_{k,n}=\alpha_{k}\exp{(-\beta_{k}n)}$ is used, and is fit to empirical data. While this captures the exponential dependence on $n$, and is thus consistent with our results, we emphasize that in our case this scaling dependence is analytically derived, and not empirically postulated. Finally, Ref.[20] studied approximability of functions with matrix product states. Assuming extendibility to the complex plane they bounded the maximal deviation between the function and its approximation with MPS with $k$-th bond dimension truncated to $\chi$ by $O(2^{-k\chi})$. While Ref.[20] does not study the entanglement properties of the encoding, this result agrees with our predictions, as the maximal deviation should be of the same order as the maximal truncated entanglement spectra, which in this case is precisely of the order $O(2^{-k\chi})$. Moreover, our results allow to compute the exact coefficients in the obtained expansions and bounds, which we demonstrate for probability distributions.
@@
III.2 Encoding probability distributions
@@
Finally, in Figs.2c,d we compute and compare the third entanglement eigenvalue $\Lambda_{k,2}$ for Lévy and normal distributions, respectively. As expected, our analytical result matches the numerical data, which indeed behave as $O(1/4^{k})$. As in the previous cases, a discrepancy with analytic results appears only on the first bonds corresponding to the large scales.
@@
III.3 Encoding dependence on the localization and support of the function
@@
The above argument is borne out by numerical experiments. In Figs.3a,b we plot the logarithm of the second entanglement eigenvalue $\Lambda_{k,1}$ on bond $k$ between qubits $k$ and $k+1$, for the normal and Lévy distributions, respectively. The large-$k$ tails are universal and exponentially decaying, as derived in Sec.III.1. The low-$k$ behaviour, however, shows indeed a super-exponential decay. Moreover, there is a clear distinction between the exponential localization of the normal distribution and the weaker polynomial one of the Lévy distribution: for the former, in Fig.3a, the large scale (small $k$) drop of entanglement is almost instantaneous, and also the universal asymptotic regime of exponential decay for lower scales (large $k$) is clearly visible already for a heavily truncated function support of $[0,2^{4}]$. Note that the maximum is achieved between $k=3,4$, consistent with the condition $k\approx\log_{2}{(L/2\sigma)}$ derived in Sec.III.1. For the Lévy case the small-$k$ decay of entanglement is much slower (though still super-exponential), and correspondingly the onset of the universal regime of entanglement decay at small scales is pushed to much larger values of $k$: we required a support truncated to a very large interval $[0,2^{19}]$ in Fig.3b to clearly demonstrate both the low- and high-$k$ regimes.
@@
IV Constructing the encoding circuit using the MPS representation
@@
V Numerical tests of the encoding circuits on simulators and QPUs
@@
In Fig. 6a we show the infidelity of a 1-layer approximating quantum circuit as a function of the truncated support size $L$, assuming a fixed discretization size (thus the qubit number is $N=10+\lceil\log_{2}L\rceil$). Here we consider the example of a normal distribution with $\sigma=2$. The infidelity is computed on an MPS simulator with a very large bond dimension (i.e. essentially exact simulation). Overall, the infidelity is excellent already for this extremely shallow circuit, no higher than $10^{-3}$. Note also that for smaller $L$ it perfectly follows the analytical estimate from Eq.(17). For larger supports this simple approximation is not accurate: as discussed in Sec.III.1 the smoothness-controlled regime is valid from bond $k=\log_{2}(L/2\sigma)$ upwards, and below the entanglement is localization controlled (and, for this exponentially localized function, much smaller, see Fig.3a). For large $L$ thus Eq.(17) overestimates the infidelity, which can be corrected by only summing contributions from bonds $k\geq log_{2}(L/2\sigma)$, i.e. the ones for which the asymptotic scaling is valid. This corrected estimate agrees perfectly with the numerical results for all $L$.
@@
In Fig.5 we examine the influence of using multiple disentangling V-layers in the quantum circuit construction, and the sensitivity to the V-layer origin position, on the example of the Lévy distribution. We observe that the fidelity can be systematically improved by adding additional disentangling layers (note, though, that in the presence of noise this is not true, as the increased circuit depth may result in a lower fidelity in practice). Furthermore, for more than a single layer we also see a strong dependence on the layer origin, as expected from the entanglement analysis in Sec.III.1: the effect is comparable in size to adding another disentangling layer, and hence the freedom to chose it is very important in practice. Note also that the minimal circuit depth is achieved with the origin on the middle qubits, further emphasizing the usefulness of this strategy.
@@
A more rigorous test of performance is obtained by computing statistical metrics for the samples drawn from the encoded distributions. We use the Kullback-Leibler (KL) divergence between the encoded and ideal distributions, and the Kolmogorov-Smirnov (KS) test. The results are reported in Tables 2 and 2. For these tests we use a 2-layer encoding circuit constructed by our complete algorithm as described in Sec.IV, including truncation of gates in V-layers set by the parameter $\varepsilon_{trunc}$. We report the depth and amount of CNOT gates of the circuit, which are computed in the RX, RY, RZ and CNOT basis (the depth includes all one-qubit gates). To collect the data we varied the interval $L$ and scale factors of the distributions, and we report the mean value and the standard deviation. The constructed distributions mostly pass the statistical tests with a wide margin, but we note that for very small amount of qubits KS test may occasionally fail (p-value$<0.05$) due to extreme discretization of the distribution, as the test is performed against the ideal truncated cumulative density function which is continuous. In Table 2 we set the hyperparameter $\varepsilon_{trunc}$ to $10^{-3}$, while in Table 2 to $\varepsilon_{trunc}$ to $10^{-4}$. Changing this value represents an accuracy-depth trade-off, which will be important on noisy QPUs.
@@ Acknowledgements
We thank Vladyslav Los and Oleksandra Hutor for their technical support in executing the quantum circuits on the IBM quantum hardware.
@@ A.1 Proof of Theorem 1
$p_{k}=\frac{1}{4^{k}}\sum_{\mathbf{u_{k},u^{\prime}_{k}}}$ $\int_{0}^{1}\int_{0}^{1}\left[f(\mathbf{u_{k}}+v/2^{k})f^{*}(\mathbf{u_{k}}+v^{\prime}/2^{k})\right.$
$\left.f(\mathbf{u^{\prime}_{k}}+v^{\prime}/2^{k})f^{*}(\mathbf{u^{\prime}_{k}}+v/2^{k})\right]\rm{d}v\rm{d}v^{\prime}.$ (20)
@@
$\sum_{u_{n}}\frac{f(u_{n})}{2^{k}}$ $=\int_{0}^{1}f(u)\rm{d}u+$ (21)
$+\frac{f(0)-f(1)}{2^{k+1}}+B_{2}\frac{f^{\prime}(1)-f^{\prime}(0)}{2^{2k+1}}+O(1/2^{3k}),$
@@
$p_{k}=\int_{0}^{1}\int_{0}^{1}\left(\int_{0}^{1}f(u+\frac{v}{2^{k}})f^{*}(u+\frac{v^{\prime}}{2^{k}})\rm{d}u+b_{1}(v,v^{\prime})\right)\\ \left(\int_{0}^{1}f^{*}(u^{\prime}+\frac{v}{2^{k}})f(u^{\prime}+\frac{v^{\prime}}{2^{k}})\rm{d}u^{\prime}+b_{2}(v,v^{\prime})\right)\rm{d}v\rm{d}v^{\prime},$ (22)
@@
$(h_{2,0}+h_{0,2})/6+h_{1,1}/4-(|f(1)^{2}|^{\prime}-|f^{2}(0)|^{\prime})/6=\\ =(h_{2,0}+2h_{1,1}+h_{0,2})/6-h_{1,1}/12-(|f(1)^{2}|^{\prime}-|f^{2}(0)|^{\prime})/6\\ =-h_{1,1}/12$ (27)